OpenStudy (anonymous):

simplify 1 + i divided by 1 - i please!!!

OpenStudy (muhammad_nauman_umair):

here is imaginary number or not?

OpenStudy (across):

\[\frac{1+i}{1-i}=\frac{(1+i)(1+i)}{(1-i)(1+i)}=i\]

OpenStudy (anonymous):

yea theres an imaginary number, and thanks!!

OpenStudy (muhammad_nauman_umair):